OpenStudy (anonymous):

question in the comments volume=1256m(cubed) diameter=? my answer: 96.6

OpenStudy (anonymous):



OpenStudy (anonymous):

was my answer correct

OpenStudy (anonymous):

is my answer correct

OpenStudy (phi):

volume is area of the base times the height the height is 13 m the area of the base is pi r^2 so pi r^2 13 = 1256 or pi r^2 = 1256/13 = 96.62 and r^2 = 96.62/pi or r^2= 30.75 (roughly) r is sqrt(30.75) = 5.55 2r is the diameter = 11.1

OpenStudy (phi):

You found the area of the base of the cylinder.

OpenStudy (anonymous):

in simpler words i get how volume is area of the base times the height, but how did you get the diameter

OpenStudy (phi):

If we call the area of the base A, we can say the volume is A*13 which equals 1256 (because they told us the volume is 1256) if we divide both sides by 13: A*13/13 = 1256/13 or A = 1256/13 or A= 96.6 (roughly) the area of the base is 96.6 square meters the base is a circle, with an unknown radius (call it r) and an unknown diameter = 2*r however, we know the area of a circle is pi*r*r in other words pi*r^2 = 96.6 divide both sides by pi r^2 = 96.6/pi = 38.75 take the square root of both sides to "solve for r" r= 5.55 meters the diameter is 2 times the radius, so the diameter is 11.1